For a certain horse race comma the odds in favor of a certain horse finishing in second place horse race, the odds in favor of a certain horse finishing in second place are given as 41 to 59

Express the indicated degree of likelihood as a probability value between 0 and 1 inclusive.

The probability is _. (Round to two decimal places as needed.)
Answer #1

since probability of an even   =odds/(1+odds)

therefore probability of horse finishing in second place =(41/59)/(1+41/59) = 0.41